A high school band held a bake sale. The number of cupcake sold was four more then twice the number of cookies sold.the band sold a total of 52cupcake and cookies. How many cupcake were sold
Answer by macston(5194)

x=number of cookies; y=number of cupcakes=2x+4
x+y=52 Substitute for y.
x+2x+4=52
3x=48
x=16 They sold 16 cookies
y=2x+4=2(16)+4=32+4=36 ANSWER: They sold 36 cupcakes
CHECK:
x+y=52
16+36=52
52=52